The difference between two numbers is 8 and twice the sum of their reciprocals are equal to 3 times the difference of their reciprocals. Find the numbers.
Answer by amarjeeth123(574)

Let one number be x.
Then the other number is (x+8).
Twice the sum of their reciprocals are equal to 3 times the difference of their reciprocals.
2(1/x+1/(x+8))=3(1/x-1/(x+8))
Simplifying we get
2((x+8)+x)=3((x+8)-x)
2(2x+8)=24
4(x+4)=24
x+4=6
x=2
The numbers are 2 and 10.